# Theoretical analysis of performance limitation of computational refocusing in optical coherence tomography

### v1

@@
High-numerical-aperture optical coherence tomography (OCT) enables sub-cellular imaging but faces a trade-off between lateral resolution and depth of focus. Computational refocusing can correct defocus in Fourier-domain OCT, yet its limitations remain unaddressed theoretically. We model the pupil-based lateral imaging process to analyze defocus and the constraints of computational refocusing in point-scanning OCT and spatially-coherent full-field OCT (FFOCT). The maximum correctable defocus (MCD) is primarily limited by confocality in point-scanning OCT, while spatially-coherent FFOCT has no such constraint, achieving virtually infinite MCD. This makes spatially-coherent FFOCT particularly suitable for optical coherence microscopy.
@@ 1 Introduction
Simultaneous cellular-level resolution and millimeter imaging depths are essential for recent applications of non-invasive tissue imaging, including tumor spheroid assessment[1], organoid visualization[2], and skin imaging[3]. The Optical coherence tomography (OCT) microscope has been demonstrated as a suitable high-resolution imaging modality for these sample types[4]. However, like other optical microscopic modality types, OCT suffers from a trade-off between its lateral resolution and its depth-of-focus (DOF). For example, a high-numerical-aperture (high-NA) objective can provide the microscope with high lateral resolution, but its DOF becomes shallow. As a result of the progress in cultivation technology, in vitro samples have come to emulate living tissues more realistically, and thus these samples are becoming thicker. The resolution-DOF trade-off has therefore become one of the important emerging problem of OCT microscopes.
@@
A short DOF is not a major issue for time-domain OCT microscope, because it measures each depth at each image acquisition. Specifically, we can introduce a technique known as “dynamic focus”[5, 6, 7], which shifts the depth position of the focus to the optimal depth for acquisition of each individual image. On the other hand, Fourier-domain (FD-) OCT with a higher NA objective, i.e., OCT microscope, is not free from the resolution-DOF trade-off.
@@
It is known that this trade-off can be mitigated using computational refocusing methods, such as holographic signal processing or interferometric synthetic aperture microscope (ISAM) techniques. The former method manipulates the phases of the spatial frequency spectrum of complex OCT signals to remove the defocus[8, 9, 10]. The latter method (ISAM) first converts a complex OCT signal into its spatial frequency spectrum and then resamples the spectrum into an appropriate space to remove the defocus [11]. A combination of the holographic signal processing and ISAM has also been demonstrated [12].
@@
Despite the increasing importance of OCT microscopy and the computational refocusing, these computational refocusing methods are expected to have certain limitations. Fukutake et al.suggested that both holographic refocusing and ISAM cannot be perfect if NA is high [13]. In holographic refocusing, the spatial sampling density of the OCT signal limits the accessible spatial frequency component and thus may govern the maximum correctable wavefront errors, including the defocus. Specifically, the correction of larger aberrations requires higher spatial frequency information, and to capture this higher spatial frequency information, higher spatial sampling density are required. The confocality may also limit the practical correctable defocus because it causes a signal energy loss, and thus, a drop in the signal-to-noise ratio (SNR), at depths far from the depth position of the physical focus. Computational refocusing techniques, including both holographic refocusing and ISAM, cannot recover this signal energy loss. Although these limitations have been anticipated, they have not been theorized well or investigated thoroughly. In addition, the effects of thiese limiting factors and the associated affecting mechanisms vary among the different types of OCT systems, e.g., scanning OCT and full-field OCT.
@@
This work aims to establish a theoretical framework that can predict the maximum correctable defocus for multiple types of OCT, including conventional point-scanning OCT and full-field (FF-) swept-source OCT with spatially coherent illumination (known hereafter as spatially-coherent FFOCT). We start with mathematical descriptions of the image formation processes for these types of OCT. Two standard OCT types, i. e., point-scanning OCT and spatially-coherent FFOCT, are considered. Here, we use the dual pupil-based formulation for OCT imaging [14, 15, 16]. Then, the defocus in the two OCT types under study is described mathematically by extending this formulation. Finally, the criteria for the maximum correctable defocus are derived for both the point-scaning OCT and spatially-coherent FFOCT. In this work, we consider two limiting factor types. The first is the sampling-density limit. Because it is related to the Nyquist frequency required to detect and correct the phase errors induced by the defocus properly, we call it the “phase-sampling limit.” The other factor is the confocality limit. Because it is related to the signal reduction at the out-of-focus depth, we call this factor the “intensity-decay limit.” Our theory will indicate that spatially-coherent FFOCT has a virtually infinitely large maximum correctable defocus and is thus particularly suitable for OCT microscopy applications.
@@ 2.1 Two-dimensional pupil-based imaging formation theory
To derive the phase sampling limit, we begin by mathematically modeling the two-dimensional (2D) lateral imaging process of OCT. The OCT lateral imaging process is described using the concepts of the “conceptual pupil”, the “spot”,the “aperture”, and the “point spread function (PSF)”[14, 17]. For simplicity, we assume that the lateral and axial resolutions are not coupled, thus allowing us to focus on the en face plane. Note here that the 2D nature of this analysis limits its applicability to the en face signal processing-based holographic refocusing approach, and it is not applicable to ISAM. This is because ISAM is based on 3D remapping of data in the spatial frequency domain.
Because we are modeling the 2D lateral imaging process, a 2D (not 3D) pupil is used. Figure 1 presents a diagram that illustrates the conceptual pupil theory.
Figure 1: Diagram of the conceptual pupil theory.
Although the conceptual pupil is related to a physical pupil in an optical system, it does not really correspond to a physically existing pupil or to a particular plane (i.e., the pupil plane) in the optics. The conceptual pupil is a representation of illumination or collection (i.e., detection) optics in the spatial frequency domain. The extension (i.e., the size) of the conceptual pupil in the frequency domain is defined by the NA of the optics, where a larger the NA corresponds to a larger the conceptual pupil. In addition, the wavefront aberration, which includes the defocus, is represented as the phase of the conceptual pupil. Because each of illumination and collection optics has its own conceptual pupil, there are two conceptual pupils in a single OCT system, i.e., the conceptual illumination and collection pupils. Hereafter, we denote the conceptual pupil as “pupil” for simplicity.
@@ 2.2 Point-scanning OCT and spatially-coherent FFOCT
In this manuscript, two types of OCT are considered: fiber-based point-scanning OCT and spatially-coherent FFOCT. spatially-coherent FFOCT is typically implemented as a swept-source OCT with plane-wave illumination. Note that time-domain FFOCT typically uses a spatially incoherence light source, and thus the discussion in this manuscript cannot be applicable to the spatially-incoherent time-domain FFOCT. On the other hand, the point-scanning OCT system in this manuscript covers time-domain, spectral-domain, and swept-source OCT systems as far as they are single-mode-fiber-based point-scanning systems.
@@ 2.2.1 Point-scanning OCT
The illumination optics of point-scanning OCT start from the single-mode fiber tip. A diffraction-limit Gaussian beam is incident from the fiber tip, diverges as it propagates, and is then collimated by a collimator to form a collimated Gaussian beam. This Gaussian beam illuminates the objective and then forms a focused beam spot on a sample. Because the physical aperture of the objective is larger than the collimated Gaussian beam, the illumination spot becomes Gaussian. This beam spot is identical to the illumination spot ${\mathrm{S_{ill}^{ps}}}({\boldsymbol{r}};k)$, where ${\boldsymbol{r}}=(x,y)$ represents the lateral coordination in the real (i.e., physical) space and $k=2\pi/\lambda$ is the wavenumber corresponds to the representative wavelength $\lambda$ of the illumination. Hereafter, we omit $k$ for simplicity, and this omission does not cause significant inaccuracy under narrowband approximation for spatially-coherent FFOCT.
@@
While conventional time-domain FF-OCT uses incoherent flood illumination, spatially-coherent FFOCT illuminates the sample using a spatially coherent plane wave, i.e., a light field with a flat phase. This indicates that the light source should be fully spatially coherent, and this condition can be achieved when the light is incident from a single-mode fiber tip, as shown in Fig. 2(b). The light is then collimated once and converged at the back focal plane of the objective, which means that it is collimated again by the objective and thus illuminates the sample as a plane wave. Specifically, the illumination spot of the spatially-coherent FFOCT is a constant ${\mathrm{S_{ill}^{sc}}}({\boldsymbol{r}};k)=\mathrm{Constant}$.
@@
The inverse Fourier transform of the cylinder function is an Airy disk function, and thus the collection spot of the spatially-coherent FFOCT is also represented by an Airy disk function. Here, we ignore the relatively small outer rings of the Airy disk and can approximate the central lobe reasonably well using a Gaussian profile [20]. Namely, we consider a virtual Gaussian collection spot ${\mathrm{S_{col}^{sc}}}({\boldsymbol{r}};k)$ for the spatially-coherent FFOCT, which is similar to the collection spot used for point-scanning OCT. It should be noted here that this Gaussian approximation results in an tacit approximation that the collection pupil ${\mathrm{S_{col}^{sc}}}({\boldsymbol{r}};k)$ also has a Gaussian.
@@ 2.3.1 Defocused Gaussian beam in point-scanning OCT
Figure 3: Schematic diagrams of probe optics used in (a) point-scanning OCT and (b) spatially-coherent FFOCT
@@
The sample plane (i.e., the en face imaging plane) is assumed to be shifted from the focus depth by $z_{d}$. Hereafter we call $z_{d}$ as the “defocus distance.” According to Ralston et al.[16], the Gaussian spot, which represents the illumination and collection spots equally, is given by
$G({\boldsymbol{r}};z_{d})=\frac{w_{0}}{w(z_{d})}\exp\left[-\frac{{\boldsymbol{r}}^{2}}{w^{2}(z_{d})}\right]\exp\left[-i\left\{nk_{0}(\Delta z)+\frac{nk_{0}{\boldsymbol{r}}\cdot{\boldsymbol{r}}}{2R(z_{d})}-\psi(z_{d})\right\}\right],$ (3)
where $\Delta z$ is the path-length difference between the reference and probe beams and $k_{0}=\frac{2\pi}{\lambda_{0}}$ is the wave number that corresponds to the center wavelength $\lambda_{0}$. $n$ is the refractive index of the sample. $w_{0}$ is the beam waist radius of the amplitude (not the intensity) at the in-focus depth, i. e., $2w_{0}$ is the diffraction-limit $1/e$-width spot size of the amplitude, and equally, the $1/e^{2}$-width spot size of the squared intensity of the spot. More specifically, $w_{0}=4f/n\phi k_{0}$, where $f$ is the focal length of the objective and $\phi$ is the $1/e^{2}$-diameter of the probe beam incident at the objective. $w(z_{d})$ is the beam radius with a particular defocus $z_{d}$, and is given by
@@
where $z_{R}=\left(n^{2}k_{0}{w_{0}}^{2}\right)/2=\left(n^{2}{w_{0}}^{2}\pi\right)/\lambda_{0}$ is the Rayleigh length. $R(z_{d})$ is the phase curvature induced by the defocus and is defined as
@@
In addition, $\psi(z_{d})$ is the Gouy phase.
@@ 2.3.2 Full-field collected defocus in SC-FFOCT
Figure 3(b) illustrates the illumination and collection used for spatially-coherent FFOCT. Because the illumination is a plane wave, the illumination spot remains a constant, irrespective of the defocus distance $z_{d}$. (See also Section 2.2.2.) Note that this insensitivity of the illumination spot to the defocus distance can be also described with respect to the illumination pupil. In general, the defocus can be described as the phase error of the pupil. Because the illumination pupil for spatially-coherent FFOCT (${\mathrm{P_{ill}^{sc}}}$) is a delta function, any phase error only causes an constant phase offset. Because the illumination spot is the inverse Fourier transform of the illumination pupil, the illumination pupil is thus not sensitive to the defocus, with the exception of a possible constant phase offset.
@@
With the normalized defocus distance $\zeta_{d}=z_{d}/z_{R}$, substitution of $R(z_{d})$(Eq.5) into ${\varphi_{\mathrm{ps}}}$ and ${\varphi_{\mathrm{sc}}}$ means that these phase, as functions of the normalzied defocus distance, then becomes, ${\varphi_{\mathrm{ps}}}({\boldsymbol{r}};\zeta_{d})=\frac{2{\boldsymbol{r}}\cdot{\boldsymbol{r}}}{nw_{0}^{2}}\frac{1}{\zeta_{d}[1+1/\zeta_{d}^{2}]}$ and ${\varphi_{\mathrm{sc}}}({\boldsymbol{r}};\zeta_{d})=\frac{{\boldsymbol{r}}\cdot{\boldsymbol{r}}}{nw_{0}^{2}}\frac{1}{\zeta_{d}[1+1/\zeta_{d}^{2}]}$, respectively. Comparison of these equations with the same defocus distance shows that the defocus-induced phase of spatially-coherent FFOCT is two times smaller than that of point-scanning OCT. Therefore, it can be deducted that the ${\boldsymbol{r}}$-dependent phase of the PSF for spatially-coherent FFOCT is two times less sensitive to the defocus distance than that of the PSF for point-scanning OCT.

### v3

@@
High-numerical-aperture optical coherence tomography (OCT) enables sub-cellular imaging but faces a trade-off between lateral resolution and depth of focus. Computational refocusing can correct defocus in Fourier-domain OCT, yet its limitations remain unaddressed theoretically. We formulate the lateral imaging process of OCT by using pupil-based imaging theory and the constraints of the computational refocusing in point-scanning OCT and spatially-coherent full-field OCT (FFOCT) are analyzed. The constrains in lateral sampling density and the confocality are considered, and it is shown that the maximum correctable defocus (MCD) is primarily limited by confocality in point-scanning OCT, while spatially-coherent FFOCT has no such constraint and can achieve virtually infinite MCD with a proper and reasonable sampling density. This makes spatially-coherent FFOCT particularly suitable for optical coherence microscopy.
@@ 1 Introduction
As a result of the progress in cultivation technology, in vitro samples have come to emulate living tissues more realistically, and thus these samples are becoming thicker. These samples and also thick ex vivo and dissected tissues have been widely used for biology and medicine, and simultaneous cellular-level-resolution and millimeter-depth imaging is essential to investigate these samples. The optical coherence tomography (OCT) microscopy[1, 2, 3] and its variants such as dynamic OCT microscope [4, 5, 6, 7] have been demonstrated for non-invasive high-resolution imaging of in vitro samples, such as spheroids [8, 7, 9, 10, 11] and organoids [12, 13], dissected tissues [14, 15, 5, 6, 16, 17, 18], and also in vivo tissues [19, 20].
However, similar to other optical microscopic modalities, OCT suffers from a trade-off between its lateral resolution and its depth-of-focus (DOF). Namely, a high-numerical-aperture (high-NA) objective can provide high lateral resolution, but its DOF becomes shallow. As the target samples became thicker, this resolution-DOF trade-off became more the important problem of OCT microscopes.
@@
A short DOF is not a major issue for time-domain OCT microscope, because it measures each depth at each image acquisition. Specifically, we can introduce a technique known as “dynamic focus”[21, 22, 23], which shifts the depth position of the focus to the optimal depth for acquisition of each individual image. On the other hand, Fourier-domain (FD-) OCT with a high NA objective, i.e., OCT microscope, is not free from the resolution-DOF trade-off.
@@
It is known that this trade-off can be mitigated using computational refocusing methods, such as holographic focus and/or aberration correction [24, 25, 26] or interferometric synthetic aperture microscope (ISAM) [27] techniques. The former method manipulates the phases of the spatial frequency spectrum of complex OCT signals to remove the defocus. The latter method (ISAM) first converts a complex OCT signal into its spatial frequency spectrum and then resamples the spectrum into an appropriate space to remove the defocus. A combination of the holographic signal processing and ISAM has also been demonstrated [28].
@@
Despite the increasing importance of OCT microscopy and the computational refocusing, these computational refocusing methods are expected to have certain limitations. In holographic refocusing, the spatial sampling density of the OCT signal limits the accessible spatial frequency component and thus may govern the maximum correctable defocus. Specifically, larger defocus causes larger local phase-error slope, and hence, consists of higher spatial frequency information. To capture this higher spatial frequency information, higher spatial sampling density are required. The confocality may also limit the practical correctable defocus because it causes an optical energy loss, and thus, a drop in the signal-to-noise ratio (SNR), at depths far from the depth position of the physical focus. Computational refocusing techniques, including both holographic refocusing and ISAM, cannot recover this optical energy loss. Although these limitations have been anticipated, they have not been theorized well or investigated thoroughly. In addition, the effects of these limiting factors and the associated affecting mechanisms vary among the different types of OCT systems, e.g., scanning OCT and full-field OCT.
@@
This work aims to establish a theoretical framework that can predict the maximum correctable defocus for multiple types of OCT, including conventional point-scanning OCT and full-field (FF-) swept-source OCT with spatially coherent illumination (known hereafter as spatially-coherent FFOCT). We start with mathematical descriptions of the image formation processes for these types of OCT. Two standard OCT types, i. e., point-scanning OCT and spatially-coherent FFOCT, are considered. Here, we use the dual pupil-based formulation for OCT imaging [29]. Then, the defocus in the two OCT types under study is described mathematically by extending this formulation. Finally, the criteria for the maximum correctable defocus are derived for both the point-scanning OCT and spatially-coherent FFOCT. In this work, we consider two limiting factor types. The first is the sampling-density limit. Because it is related to the Nyquist frequency required to detect and correct the phase errors induced by the defocus properly, we call it the “Nyquist criterion.” The other factor is the confocality limit. Because it is related to the signal reduction at the out-of-focus depth, we call this factor the “confocality-limit criterion.” Our theory will indicate that spatially-coherent FFOCT has a virtually infinitely large maximum correctable defocus and is thus particularly suitable for OCT microscopy applications.
@@ 2.1 Two-dimensional pupil-based imaging formation theory
To derive the phase sampling limit, we start from mathematical modeling of the two-dimensional (2D) lateral imaging process of OCT. Here the imaging process is described using the concepts of the “conceptual pupil”, the “spot”,the “aperture”, and the “point spread function (PSF)”[30, 31, 32]. For simplicity, we assume that the lateral and axial resolutions are not coupled, thus allowing us to focus our consideration on the en face plane. Note that the 2D nature of this analysis limits its applicability to the en face signal processing-based holographic refocusing approach, and it is not applicable to ISAM because ISAM is based on 3D remapping of data in the spatial frequency domain.
Because we are modeling the 2D lateral imaging process, a 2D (i.e., not 3D) conceptual pupil is used. Figure 1 presents a diagram that illustrates the conceptual pupil theory.
Figure 1: Diagram illustrating the interrelationship among the conceptual pupils, spots, aperture, and point spread function. $(x,y)$ and $(k_{x},k_{y})$ are the lateral spatial coordinates and their corresponding spatial frequencies, respectively. FT stands for Fourier transform.
Although the conceptual pupil is somehow related to a physical pupil in an optical system, it does not really correspond to the physically existing pupil or to a particular plane (i.e., the pupil plane) in the optics. The conceptual pupil is a representation of illumination or collection (i.e., detection) optics in the spatial frequency domain. The extension (i.e., the size) of the conceptual pupil in the frequency domain is defined by the NA of the optics, where a larger the NA corresponds to a larger the conceptual pupil. In addition, the wavefront aberration, which includes the defocus, is represented as the phase of the conceptual pupil. Because each of illumination and collection optics has its own conceptual pupil, there are two conceptual pupils in a single OCT system, i.e., the conceptual illumination and collection pupils, as illustrated at the top left of Fig. 1. Hereafter, we denote the conceptual pupil as “pupil” for simplicity.
@@ 2.2 Point-scanning OCT and spatially-coherent FFOCT
In this manuscript, two types of OCT are considered: fiber-based point-scanning OCT and spatially-coherent FFOCT. Spatially-coherent FFOCT is typically implemented as a swept-source OCT with plane-wave illumination. Note that time-domain FFOCT typically uses a spatially incoherent light source, and thus the discussion in this manuscript cannot be applicable to the spatially-incoherent time-domain FFOCT. On the other hand, the point-scanning OCT system in this manuscript covers time-domain, spectral-domain, and swept-source OCT systems as far as they are single-mode-fiber-based point-scanning systems.
@@ 2.2.1 Point-scanning OCT
The illumination optics of point-scanning OCT start from the single-mode fiber tip. A diffraction-limit Gaussian beam is incident from the fiber tip, diverges as it propagates, and is then collimated by a collimator to form a collimated Gaussian beam. This Gaussian beam illuminates the objective and then forms a focused beam spot on a sample. Because the physical aperture of the objective is larger than the collimated Gaussian beam, the illumination spot becomes Gaussian. This beam spot is identical to the illumination spot ${\mathrm{S_{ill}^{ps}}}({\boldsymbol{r}};k)$, where ${\boldsymbol{r}}=(x,y)$ represents the lateral coordination in the real (i.e., physical) space and $k=2\pi/\lambda$ is the wavenumber corresponds to the representative wavelength $\lambda$ of the illumination. Hereafter, we omit $k$ for simplicity, and this omission does not cause significant inaccuracy under relatively narrow-band approximation.
@@ 2.2.2 Spatially-coherent FFOCT
While conventional time-domain FF-OCT uses incoherent flood illumination, spatially-coherent FFOCT illuminates the sample using a spatially coherent plane wave, i.e., a light field with a flat phase. This indicates that the light source should be fully spatially coherent, and this condition can be achieved when the light is incident from a single-mode fiber tip, as shown in Fig. 2(b). The light is then collimated once and converged at the back focal plane of the objective, which means that it is collimated again by the objective and thus illuminates the sample as a plane wave. Specifically, the illumination spot of the spatially-coherent FFOCT is a constant as ${\mathrm{S_{ill}^{sc}}}({\boldsymbol{r}};k)=\mathrm{Constant}$.
@@
The inverse Fourier transform of the cylinder function is an Airy disk function, and thus the in-focus collection spot of the spatially-coherent FFOCT is also represented by an Airy disk function. Here, we ignore the relatively small outer rings of the Airy disk and can approximate the central lobe reasonably well using a Gaussian profile [35]. The out-of-focus collection spot is assumed to be the convolution of the approximated Gaussian in-focus spot and the phase function. Namely, we consider a virtual Gaussian collection spot ${\mathrm{S_{col}^{sc}}}({\boldsymbol{r}};k)$ for the spatially-coherent FFOCT, which is similar to the collection spot used for point-scanning OCT. It should be noted here that this Gaussian approximation results in a tacit approximation that the collection pupil ${\mathrm{S_{col}^{sc}}}({\boldsymbol{r}};k)$ also has a Gaussian.
@@ 2.3.1 Defocused Gaussian beam in point-scanning OCT
Figure 3: Schematic diagrams of probe optics used in (a) point-scanning OCT and (b) spatially-coherent FFOCT.
@@
The sample plane (i.e., the en face imaging plane) is assumed to be shifted from the focus depth by $z_{d}$. Hereafter we call $z_{d}$ as the “defocus distance.” The Gaussian spot, which represents the illumination and collection spots equally, can be given by using the Gaussian beam [36] as
$G({\boldsymbol{r}};z_{d})=\frac{w_{0}}{w(z_{d})}\exp\left[-\frac{{\boldsymbol{r}}^{2}}{w^{2}(z_{d})}\right]\exp\left[-i\left\{nk_{0}(z_{d})+\frac{nk_{0}{\boldsymbol{r}}\cdot{\boldsymbol{r}}}{2R(z_{d})}-\psi(z_{d})\right\}\right],$ (3)
where $w_{0}$ is the beam waist radius of the amplitude (not the intensity) at the in-focus depth, i. e., $2w_{0}$ is the diffraction-limit $1/e$-width spot size of the amplitude, and equally, the $1/e^{2}$-width spot size of the squared intensity of the spot. More specifically, $w_{0}=4f/\phi k_{0}$, where $k_{0}=\frac{2\pi}{\lambda_{0}}$ is the wave number that corresponds to the center wavelength $\lambda_{0}$, $n$ is the refractive index of the sample, $f$ is the focal length of the objective and $\phi$ is the $1/e^{2}$-diameter of the probe beam incident at the objective. $w(z_{d})$ is the beam radius with a particular defocus $z_{d}$, and is given by
@@
where $z_{R}=\left(nk_{0}{w_{0}}^{2}\right)/2=\left(n{w_{0}}^{2}\pi\right)/\lambda_{0}$ is the Rayleigh length. The second term in the second exponential represents the quadratic phase induced by defocus and $R(z_{d})$ is the phase curvature:
@@
In the third term of the second exponential, $\psi(z_{d})$ is the Gouy phase.
@@ 2.3.2 Full-field collected defocus in SC-FFOCT
Figure 3(b) illustrates the illumination and collection used for spatially-coherent FFOCT. Because the illumination is a plane wave, the illumination spot remains a constant, irrespective of the defocus distance $z_{d}$ (see also Section 2.2.2.) Note that this insensitivity of the illumination spot to the defocus distance can be also described with respect to the illumination pupil. In general, the defocus can be described as the phase error of the pupil. Because the illumination pupil for spatially-coherent FFOCT ($\mathrm{P_{ill}^{sc}}$}) is a delta function, any phase error only causes an constant phase offset. Because the illumination spot is the inverse Fourier transform of the illumination pupil, the illumination pupil is thus not sensitive to the defocus, with the exception of a possible constant phase offset.
@@
With the normalized defocus distance $\zeta_{d}=z_{d}/z_{R}$, substitution of $R(z_{d})$ [Eq. (5)] into $\varphi_{\mathrm{ps}}$} and $\varphi_{\mathrm{sc}}$} means that these phase, as functions of the normalzied defocus distance, then becomes, ${\varphi_{\mathrm{ps}}}({\boldsymbol{r}};\zeta_{d})=\frac{2{\boldsymbol{r}}\cdot{\boldsymbol{r}}}{w_{0}^{2}}\frac{1}{\zeta_{d}[1+1/\zeta_{d}^{2}]}$ and ${\varphi_{\mathrm{sc}}}({\boldsymbol{r}};\zeta_{d})=\frac{{\boldsymbol{r}}\cdot{\boldsymbol{r}}}{w_{0}^{2}}\frac{1}{\zeta_{d}[1+1/\zeta_{d}^{2}]}$, respectively. Comparison of these equations with the same defocus distance shows that the defocus-induced phase of spatially-coherent FFOCT is two times smaller than that of point-scanning OCT. Therefore, it can be deducted that the $\boldsymbol{r}$}-dependent phase of the PSF for spatially-coherent FFOCT is two times less sensitive to the defocus distance than that of the PSF for point-scanning OCT.